OpenStudy (anonymous):

Suppose f (x) = x^-1 on the interval [-1, 2]. Use the Mean Value Theorem, if it applies, to find all values c in the open interval (-1, 2) such that f'(c) = (f(2) - f(-1))/(2-(-1)). Options: c= -1/2 c= 1/2 c= 9/2 Mean Value Theorem does not apply

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

your function is not defined on all real numbers is it?

OpenStudy (misty1212):

what is the domain of \[f(x)=\frac{1}{x}\]?

OpenStudy (anonymous):

Hi! :) That was the entire question, I'm not sure what the domain of f(x) = 1/x is :o

OpenStudy (misty1212):

i don't believe you

OpenStudy (misty1212):

i mean i believe it was the entire question, i don't believe you do not know where \(\frac{1}{x}\) is undefined

OpenStudy (anonymous):

Ohhh sorry I didn't understand what you were asking :) That would be when x=0

OpenStudy (misty1212):

right it is not defined at \(x=0\)

OpenStudy (misty1212):

so if you read the hypothesis of the mvt you will see that the function has to be continuous on the closed interval \([a,b]\) and differentiable in the open interval \((a,b)\)

OpenStudy (misty1212):

is your function continuous on \([-1,2]\)?

OpenStudy (anonymous):

No it would be closed :)

OpenStudy (misty1212):

??

OpenStudy (misty1212):

"closed" is a property of the interval, not of the function

OpenStudy (anonymous):

Er sorry it wouldn't be continuous

OpenStudy (misty1212):

the question is, is \(f(x)=\frac{1}{x}\) continuous on the entire interval \([1,2]\)?

OpenStudy (misty1212):

right, it is not

OpenStudy (misty1212):

because it is not even defined at 0, so it certainly cannot be continuous there, let alone differentiable

OpenStudy (misty1212):

therefore the mean value theorem says nothing in this case, because it does not fit the hypothesis

OpenStudy (anonymous):

Ohhhhh that makes so much more sense! Thank you very much! :)

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]

OpenStudy (anonymous):

Have a great day and a Happy New Year! :)

OpenStudy (xapproachesinfinity):



OpenStudy (xapproachesinfinity):

this picture is worth 1000 word :)